# SignX: Continuous Sign Recognition in Compact Pose-Rich Latent Space

### v2

@@ 5.2.1 Ablation Results
GEU 9556136 49.9 - SignBERT hu2021signbert 39.40
SLT SLT 24.5 32.0 BEST zhao2023best 46.25
Bert 10542663 21.2 30.8 SignBERT+ hu2023signbert+ 48.85
Bert fine-tuning 10542663 20.4 30.0 MSLU (zhou2024scaling) 56.29
CorrNet hu2023continuous 20.5 30.1 NLA-SLR zuo2023natural 61.05
GPGN 10542663 20.5 30.0 Uni-Sign li2025uni 63.52
CorrNet+ hu2024corrnetsignlanguagerecognition 19.1 28.2 Sigma pu2025sigmasemanticallyinformativepretraining 64.40
@@
Table 4: Comprehensive Comparison on Continuous and Isolated SLR. Left: WER on RWTH-2014-T and CSL-Daily. Right: P-I accuracy on WLASL-2000. Our SignX framework achieves SOTA performance across all tracks, even when compared with the latest Signma pu2025sigmasemanticallyinformativepretraining from the same period.
@@ 5.3 Comparison with Previous Work
We evaluate the SignX framework across multiple mainstream benchmarks 8578910; Zhou2021ImprovingSLT-with-monolingual-CSLDaily; li2020word, utilizing a comprehensive suite of metrics including WER, BLEU, and P-I accuracy. As illustrated in tables˜3 and 4, SignX achieves significant improvements over current state-of-the-art methods in both continuous sign recognition and isolated sign recognition tasks (Note that some of the SLT work is different from ours and cannot be directly compared, e.g., zhang2025largesignlanguagemodels), our framework establishes a new performance ceiling, exhibiting superior robustness and cross-domain scalability, thereby providing a high-performance baseline for CSLR task.
@@ 6 Conclusion
This paper presents SignX, a high-efficiency framework for CSLR. (i) By leveraging Vision Transformers (ViT) dosovitskiy2020vit, we construct a unified and information-dense latent space that successfully integrates heterogeneous pose representations from five state-of-the-art estimators. (ii) This latent space is further refined through multi-faceted techniques for compression, structural organization, and semantic alignment, establishing a robust sign-specific representation space that effectively maps raw sign language videos to latent embeddings. (iii) Furthermore, we perform continuous recognition within this compressed, pose-rich space by employing a multi-stage modeling approach that combines ResNet-driven zhang2023sltunet temporal feature distillation with Transformer-based sequence refinement, supplemented by latent optimization strategies such as covariance whitening and adaptive pruning. (iv) SignX significantly enhances recognition accuracy while drastically reducing computational overhead. Extensive experimental validation across four major benchmarks —demonstrates that SignX achieves SOTA performance across all metrics.
@@
Dataset Duration(h) Vocabulary(k) Annotation Type Year Domain
train val test train val test
KETI Ko-2019-SLT-based-human-keypoint-estimation 20.05 2.24 5.70 $\leftarrow$ 0.49 $\rightarrow$ Spoken Text 2019 Emergency situations
PHOENIX-2014T 2018_Neural_Sign_Language_Translation 9.2 0.6 0.7 2 0.9 1 Spoken Text, Gloss 2018 Weather Forecast
CSL Daily Zhou2021ImprovingSLT-with-monolingual-CSLDaily 20.62 1.24 1.41 2 1.3 1.3 Spoken Text, Gloss 2021 Daily life
OpenASL shi-etal-2022-openASL $\leftarrow$ 288 $\rightarrow$ $\leftarrow$ 33 $\rightarrow$ Spoken Text 2022 Youtube (news + vlogs)
How2Sign Duarte_2021_how2sign 69.6 3.9 5.6 15.6 3.2 3.6 Spoken Text 2021 Instructional
ASLLRP neidle2022asl $\leftarrow$ 80 $\rightarrow$ $\leftarrow$ 2.1 $\rightarrow$ Gloss 2022 Comprehensive
Table 5: Comparison of sign language datasets by annotation type and task focus: The top five rows primarily focus on translation to spoken word text. However, our main focus is on recognition of ASL signs (identified by ID Gloss labels) from poses and sign language videos. $\leftarrow$ $\rightarrow$ indicate that in some cases only statistics on the whole dataset are provided.
@@ A.1 Sign Language Recognition
Sign language recognition is an important research direction, at the intersection of computer vision and natural language processing. Early works mainly relied on hand-crafted features and rules to recognize small sign language vocabularies grobel1997isolated; kadir2004minimal. Subsequently, methods based on the use of machine learning (HMMs) and 3D human pose and motion analytics were used to recognize ASL 710744; 799236. With the development of deep learning, neural network-based methods began to dominate this field. The earliest deep learning methods used CNN hershey2017cnn to extract spatial features from video frames and then used RNN seq2seq-RNN to model temporal relationships cui2017recurrent; koller2015continuous. Although these methods were simple and direct, they struggled to capture fine-grained dynamic movements fang2025stablesignerhierarchicalsign.
Subsequently, researchers began to introduce human pose estimation technology Katz-IEEE-1987-estimation; Ko-2019-SLT-based-human-keypoint-estimation; charles2014automatic; SLT; SLT_pose_amit; yin2022mlslt, converting sign language videos into skeleton sequences for recognition, which significantly improved the models’ ability to perceive action details saunders2020progressive; saunders2021continuous. With the advancement of computer vision technology, various powerful pose estimation methods cai2023smplerx; yang2023effective have emerged. However, because of their different data formats and representation methods, finding a way to effectively integrate this heterogeneous information has become an urgent challenge stoll2020text2sign; huang2021towards; fang2025signllmsignlanguageproduction; fang2025signdiffdiffusionmodelamerican. As shown in Table 5, different datasets focus on different output goals. Understanding also requires taking account of the complex syntax of these languages. This involves not only linear sign order, but also interactions with non-manual signals, which play an essential grammatical role in signed languages. Gloss-based representations like “I-GO-STORE” can help, in part, to model the unique linguistic rules of a given signed language camgoz2018neural; camgoz2020sign; gloss; WordsAreOurGlosses; zelinka2020neural; neidle2022alternative; zhou-etal-2024-multimodal; neidle2022aslvideocorpora; zhang2025largesignlanguagemodels.
@@ A.2 Pose Estimation
Pose estimation technology has made significant progress in recent years. SMPLer-X cai2023smplerx achieves precise modeling of the body, hands, and face by providing complete 3D human body model parameter estimation smpl; cao2018openpose. DWPose yang2023effective focuses on 2D keypoint detection, achieving real-time performance while ensuring stability wang2018video; MediaPipe. Google’s MediaPipe MediaPipe provides lightweight but accurate 3D pose prediction, making real-time applications possible MediaPipe; chan2019everybody.
Beyond traditional pose estimation, there are also emerging technologies. PrimeDepth zavadski2024primedepth, specifically designed for depth estimation, can provide 3D structural information about scenes qiu2021dense; carreira-2017-i3d. Sapiens segmentation khirodkar2024sapiens provides a new perspective for interpreting signers’ fine-grained actions through precise human body part segmentation guler2018densepose; saunders2021mixed. These advances provide rich feature representation methods for SL recognition, offering good extensions to the information dimensions of pose estimation 7298594; koller2020quantitative.
@@ A.3 Diffusion Models
Diffusion models have achieved breakthroughs in the generative arena in recent years. From the initial DDPM ho2020denoising to Stable Diffusion rombach2021highresolution, diffusion models have demonstrated powerful image generation capabilities saunders2021skeletal; wang2018high. Researchers have also begun applying diffusion to discrete text sequence generation, with promising results ko2019neural; Bohacek_2022_WACV. Particularly in cross-modal conversion tasks, diffusion models have shown excellent performance, providing important references for converting SL videos to text saunders2020progressive; cooper2007large; zhou-etal-2024-multimodal.
The success of the stable diffusion model lies in the fact that it conducts learning not at the pixel level but in the latent space fang2025streamflowtheoryalgorithmimplementation. This enables its learning to be not only smooth but also with a reduced computational burden. Inspired by its success, our work also to apply sign language recognition in the sign latent space, and as a result, we have achieved very good results.
@@ A.4 Vision Transformer
Since the development of ViT dosovitskiy2020vit, the use of Transformers in vision tasks has become increasingly widespread. Compared to traditional architectures, Transformers can better model long-range dependencies through self-attention mechanisms saunders2021continuous; huang2021towards. The original ViT first applied pure Transformer structure to image classification, pioneering a new paradigm of vision Transformers. Subsequent improvements, such as Swin Transformer and other models, further enhanced model performance in vision tasks by introducing local attention mechanisms camgoz2020sign; ko2019neural. In the field of video event recognition, researchers have introduced spatiotemporal attention mechanisms, enabling ViT to better process video sequence data Bohacek_2022_WACV; forster2012rwth. This advance is particularly important for SL recognition, as SL videos contain rich temporal information. Through well-designed spatiotemporal attention mechanisms, models can effectively capture temporal dependencies in sequences of signing koller15:cslr; saunders2021mixed.
@@ Appendix B Methodology for Constructing Sign Language Latent Space
Our process of constructing the rich pose latent space utilize a linguistically annotated ASL dataset of continuous signing: the ASLLRP SignStream® 3 Corpus neidle2022asl. This dataset contains over 80 hours of American Sign Language (ASL) videos with synchronized front view, side view, and facial close-up recordings. These videos have been linguistically annotated using SignStream® neidle2018 software (see https://www.bu.edu/asllrp/SignStream/3/), providing detailed manually annotated gloss labels, sign types, handshape information, and non-manual grammatical markers. The dataset comprises 2,127 utterances with 17,522 sign tokens, plus an additional set of 542 sentences from DawnSignPress.
For video processing, we apply five pose extraction models: SMPLer-X cai2023smplerx for accurate 3D human body parameters, DWPose yang2023effective for real-time 2D keypoint detection, MediaPipe MediaPipe for lightweight 3D pose prediction, PrimeDepth zavadski2024primedepth for hand depth information, and Sapiens segmentation khirodkar2024sapiens for fine-grained body part segmentation. Unlike previous approaches, our preprocessing strategy retains some incomplete or slightly occluded video samples to enhance model robustness to quality variations in real-world scenarios. Videos are standardized to 30fps and processed at uniform spatial resolution. We extract multimodal pose features from all videos to form a unified representation space.
@@ B.1 Stage 1: ViT Pose2Gloss Part Training
Following seq2seq principles sutskever2014sequence, we employ a cross-entropy text generation loss mao2023cross that measures the probabilistic divergence between predicted and ground truth token (word index in codebook) sequences:
@@ B.2 Stage 2: ViT Video2Pose Part Training
Following feature transfer learning principles yosinski2014transferable, we freeze the Pose2Gloss module and train the Video2Pose component by minimizing the reconstruction error:
@@ B.3 Implementation Details of Latent Space
For Stages 1 and 2, we use the linguistically annotated ASLLRP SignStream® 3 Corpus of continuous signing neidle2022asl as primary datasets for American Sign Language (ASL) recognition. During Stage 1, we initially found that the heterogeneity and varying formats of pose features from diverse sources had a significant negative impact on the structural integrity of our latent space. To address this, we prioritized optimizing the cross-modal alignment and feature consistency in the Pose2Gloss module (Stage 1) through a gloss-aware latent distillation loss. However, we observed that relying solely on feature-level distillation did not sufficiently address the nuances of sign-to-text generation quality.
@@ E.2 Rigorous Benchmarking on the ASLLRP Dataset
As ASLLRP is a relatively new dataset for the CSLR task, there is a lack of publicly available baseline results from modern models. To ensure a fair comparison under identical pose input conditions, we re-implemented and fine-tuned several representative state-of-the-art (SOTA) frameworks, including SLTUNET zhang2023sltunet and C2SLR zuo2022c2slr. Our selection criteria prioritized mainstream methods that represent established technical trajectories in sign language understanding, rather than solely focusing on the most recent publication dates.
@@ E.3 Component Customization and Scalability
All architectural components in SignX have been deeply modified to serve the objective of latent space recognition. For instance, our Custom CodeBook is utilized to learn and initialize pose-rich features beneficial for linguistic mapping. This design is inherently scalable to an unlimited vocabulary and does not restrict the model’s capacity for future expansion. Furthermore, while our framework utilizes a comprehensive set of modalities, these modules are highly reusable across different datasets. Compared to ultra-large-scale models such as Uni-Sign li2025uni, our approach achieves superior performance with a smaller data footprint, demonstrating a data-frugal and efficient research path.
@@ E.5 Discussion on the overall motivation and methodology
Overall, our key insight is that not only should we perform generation tasks in the Latent Space, but we should also perform recognition tasks in the Latent Space. This was deemed successful by the experiments. Initially, we had the option of using pre-trained components such as VAE kingma2013auto; dong2022act and DINO3 simeoni2025dinov3 as alternatives, but they lacked pose-level information, lacked effective sign language semantic information, and the simple CNN in VAE could not meet our needs for scalability and stability. Therefore, after careful consideration, we adopted the existing methods. In the future, we will make more resources available, such as code and preprocessed data, to support the sign language research community.

### v3

@@ 5.2.1 Ablation Results
GEU Tang et al. (2022) 49.9 - SignBERT Hu et al. (2021) 39.40
SLT SLT 24.5 32.0 BEST Zhao et al. (2023) 46.25
Bert Guo et al. (2024) 21.2 30.8 SignBERT+ Hu et al. (2023a) 48.85
Bert fine-tuning Guo et al. (2024) 20.4 30.0 MSLU (Zhou et al., 2025) 56.29
CorrNet Hu et al. (2023b) 20.5 30.1 NLA-SLR Zuo et al. (2023) 61.05
GPGN Guo et al. (2024) 20.5 30.0 Uni-Sign Li et al. (2025) 63.52
CorrNet+ Hu et al. (2024) 19.1 28.2 Sigma Pu et al. (2025) 64.40
@@
Table 4: Comprehensive Comparison on Continuous and Isolated SLR. Left: WER on RWTH-2014-T and CSL-Daily. Right: P-I accuracy on WLASL-2000. Our SignX framework achieves SOTA performance across all tracks, even when compared with the latest Signma Pu et al. (2025) from the same period.
@@ 5.3 Comparison with Previous Work
We evaluate the SignX framework across multiple mainstream benchmarks Camgoz et al. (2018); Zhou2021ImprovingSLT-with-monolingual-CSLDaily; Li et al. (2020), utilizing a comprehensive suite of metrics including WER, BLEU, and P-I accuracy. As illustrated in tables˜3 and 4, SignX achieves significant improvements over current state-of-the-art methods in both continuous sign recognition and isolated sign recognition tasks (Note that some of the SLT work is different from ours and cannot be directly compared, e.g., Zhang et al. (2025)), our framework establishes a new performance ceiling, exhibiting superior robustness and cross-domain scalability, thereby providing a high-performance baseline for CSLR task.
@@ 6 Conclusion
This paper presents SignX, a high-efficiency framework for CSLR. (i) By leveraging Vision Transformers (ViT) Dosovitskiy et al. (2021), we construct a unified and information-dense latent space that successfully integrates heterogeneous pose representations from five state-of-the-art estimators. (ii) This latent space is further refined through multi-faceted techniques for compression, structural organization, and semantic alignment, establishing a robust sign-specific representation space that effectively maps raw sign language videos to latent embeddings. (iii) Furthermore, we perform continuous recognition within this compressed, pose-rich space by employing a multi-stage modeling approach that combines ResNet-driven Zhang et al. (2023) temporal feature distillation with Transformer-based sequence refinement, supplemented by latent optimization strategies such as covariance whitening and adaptive pruning. (iv) SignX significantly enhances recognition accuracy while drastically reducing computational overhead. Extensive experimental validation across four major benchmarks —demonstrates that SignX achieves SOTA performance across all metrics.
@@ A.1 Sign Language Recognition
Sign language recognition is an important research direction, at the intersection of computer vision and natural language processing. Early works mainly relied on hand-crafted features and rules to recognize small sign language vocabularies grobel1997isolated; kadir2004minimal. Subsequently, methods based on the use of machine learning (HMMs) and 3D human pose and motion analytics were used to recognize ASL Vogler and Metaxas (1998); Metaxas (1999). With the development of deep learning, neural network-based methods began to dominate this field. The earliest deep learning methods used CNN hershey2017cnn to extract spatial features from video frames and then used RNN seq2seq-RNN to model temporal relationships cui2017recurrent; koller2015continuous. Although these methods were simple and direct, they struggled to capture fine-grained dynamic movements Fang et al. (2025b).
Subsequently, researchers began to introduce human pose estimation technology Katz-IEEE-1987-estimation; Ko-2019-SLT-based-human-keypoint-estimation; charles2014automatic; SLT; SLT_pose_amit; Yin et al. (2022), converting sign language videos into skeleton sequences for recognition, which significantly improved the models’ ability to perceive action details saunders2020progressive; saunders2021continuous. With the advancement of computer vision technology, various powerful pose estimation methods Cai et al. (2023); Yang et al. (2023) have emerged. However, because of their different data formats and representation methods, finding a way to effectively integrate this heterogeneous information has become an urgent challenge stoll2020text2sign; huang2021towards; Fang et al. (2025a, c). As shown in Table 5, different datasets focus on different output goals. Understanding also requires taking account of the complex syntax of these languages. This involves not only linear sign order, but also interactions with non-manual signals, which play an essential grammatical role in signed languages. Gloss-based representations like “I-GO-STORE” can help, in part, to model the unique linguistic rules of a given signed language camgoz2018neural; camgoz2020sign; gloss; WordsAreOurGlosses; zelinka2020neural; Neidle and Ballard (2022); Zhou et al. (2024); Neidle et al. (2022b); Zhang et al. (2025).
@@ A.2 Pose Estimation
Pose estimation technology has made significant progress in recent years. SMPLer-X Cai et al. (2023) achieves precise modeling of the body, hands, and face by providing complete 3D human body model parameter estimation Loper et al. (2015); cao2018openpose. DWPose Yang et al. (2023) focuses on 2D keypoint detection, achieving real-time performance while ensuring stability wang2018video; MediaPipe. Google’s MediaPipe MediaPipe provides lightweight but accurate 3D pose prediction, making real-time applications possible MediaPipe; chan2019everybody.
Beyond traditional pose estimation, there are also emerging technologies. PrimeDepth Zavadski et al. (2024), specifically designed for depth estimation, can provide 3D structural information about scenes qiu2021dense; carreira-2017-i3d. Sapiens segmentation Khirodkar et al. (2024) provides a new perspective for interpreting signers’ fine-grained actions through precise human body part segmentation Güler et al. (2018); saunders2021mixed. These advances provide rich feature representation methods for SL recognition, offering good extensions to the information dimensions of pose estimation Szegedy et al. (2015); koller2020quantitative.
@@ A.3 Diffusion Models
Diffusion models have achieved breakthroughs in the generative arena in recent years. From the initial DDPM Ho et al. (2020) to Stable Diffusion Rombach et al. (2021), diffusion models have demonstrated powerful image generation capabilities saunders2021skeletal; wang2018high. Researchers have also begun applying diffusion to discrete text sequence generation, with promising results ko2019neural; Boháček and Hrúz (2022). Particularly in cross-modal conversion tasks, diffusion models have shown excellent performance, providing important references for converting SL videos to text saunders2020progressive; cooper2007large; Zhou et al. (2024).
The success of the stable diffusion model lies in the fact that it conducts learning not at the pixel level but in the latent space Fang et al. (2025d). This enables its learning to be not only smooth but also with a reduced computational burden. Inspired by its success, our work also to apply sign language recognition in the sign latent space, and as a result, we have achieved very good results.
@@ A.4 Vision Transformer
Since the development of ViT Dosovitskiy et al. (2021), the use of Transformers in vision tasks has become increasingly widespread. Compared to traditional architectures, Transformers can better model long-range dependencies through self-attention mechanisms saunders2021continuous; huang2021towards. The original ViT first applied pure Transformer structure to image classification, pioneering a new paradigm of vision Transformers. Subsequent improvements, such as Swin Transformer and other models, further enhanced model performance in vision tasks by introducing local attention mechanisms camgoz2020sign; ko2019neural. In the field of video event recognition, researchers have introduced spatiotemporal attention mechanisms, enabling ViT to better process video sequence data Boháček and Hrúz (2022); forster2012rwth. This advance is particularly important for SL recognition, as SL videos contain rich temporal information. Through well-designed spatiotemporal attention mechanisms, models can effectively capture temporal dependencies in sequences of signing Koller et al. (2015); saunders2021mixed.
@@ Appendix B Methodology for Constructing Sign Language Latent Space
Our process of constructing the rich pose latent space utilize a linguistically annotated ASL dataset of continuous signing: the ASLLRP SignStream® 3 Corpus Neidle et al. (2022a). This dataset contains over 80 hours of American Sign Language (ASL) videos with synchronized front view, side view, and facial close-up recordings. These videos have been linguistically annotated using SignStream® Neidle et al. software (see https://www.bu.edu/asllrp/SignStream/3/), providing detailed manually annotated gloss labels, sign types, handshape information, and non-manual grammatical markers. The dataset comprises 2,127 utterances with 17,522 sign tokens, plus an additional set of 542 sentences from DawnSignPress.
For video processing, we apply five pose extraction models: SMPLer-X Cai et al. (2023) for accurate 3D human body parameters, DWPose Yang et al. (2023) for real-time 2D keypoint detection, MediaPipe MediaPipe for lightweight 3D pose prediction, PrimeDepth Zavadski et al. (2024) for hand depth information, and Sapiens segmentation Khirodkar et al. (2024) for fine-grained body part segmentation. Unlike previous approaches, our preprocessing strategy retains some incomplete or slightly occluded video samples to enhance model robustness to quality variations in real-world scenarios. Videos are standardized to 30fps and processed at uniform spatial resolution. We extract multimodal pose features from all videos to form a unified representation space.
@@ B.1 Stage 1: ViT Pose2Gloss Part Training
Following seq2seq principles Sutskever et al. (2014), we employ a cross-entropy text generation loss Mao et al. (2023) that measures the probabilistic divergence between predicted and ground truth token (word index in codebook) sequences:
@@ B.2 Stage 2: ViT Video2Pose Part Training
Following feature transfer learning principles Yosinski et al. (2014), we freeze the Pose2Gloss module and train the Video2Pose component by minimizing the reconstruction error:
@@ B.3 Implementation Details of Latent Space
For Stages 1 and 2, we use the linguistically annotated ASLLRP SignStream® 3 Corpus of continuous signing Neidle et al. (2022a) as primary datasets for American Sign Language (ASL) recognition. During Stage 1, we initially found that the heterogeneity and varying formats of pose features from diverse sources had a significant negative impact on the structural integrity of our latent space. To address this, we prioritized optimizing the cross-modal alignment and feature consistency in the Pose2Gloss module (Stage 1) through a gloss-aware latent distillation loss. However, we observed that relying solely on feature-level distillation did not sufficiently address the nuances of sign-to-text generation quality.
@@ E.2 Rigorous Benchmarking on the ASLLRP Dataset
As ASLLRP is a relatively new dataset for the CSLR task, there is a lack of publicly available baseline results from modern models. To ensure a fair comparison under identical pose input conditions, we re-implemented and fine-tuned several representative state-of-the-art (SOTA) frameworks, including SLTUNET Zhang et al. (2023) and C2SLR Zuo and Mak (2022). Our selection criteria prioritized mainstream methods that represent established technical trajectories in sign language understanding, rather than solely focusing on the most recent publication dates.
@@ E.3 Component Customization and Scalability
All architectural components in SignX have been deeply modified to serve the objective of latent space recognition. For instance, our Custom CodeBook is utilized to learn and initialize pose-rich features beneficial for linguistic mapping. This design is inherently scalable to an unlimited vocabulary and does not restrict the model’s capacity for future expansion. Furthermore, while our framework utilizes a comprehensive set of modalities, these modules are highly reusable across different datasets. Compared to ultra-large-scale models such as Uni-Sign Li et al. (2025), our approach achieves superior performance with a smaller data footprint, demonstrating a data-frugal and efficient research path.
@@ E.5 Discussion on the overall motivation and methodology
Overall, our key insight is that not only should we perform generation tasks in the Latent Space, but we should also perform recognition tasks in the Latent Space. This was deemed successful by the experiments. Initially, we had the option of using pre-trained components such as VAE Kingma and Welling (2013); dong2022act and DINO3 Siméoni et al. (2025) as alternatives, but they lacked pose-level information, lacked effective sign language semantic information, and the simple CNN in VAE could not meet our needs for scalability and stability. Therefore, after careful consideration, we adopted the existing methods. In the future, we will make more resources available, such as code and preprocessed data, to support the sign language research community.